# How to Find Fantastic Papers: Self-Rankings as a Powerful Predictor of Scientific Impact Beyond Peer Review

### v1

@@ 1 Introduction
A central objective of peer review is to identify and promote research of high impact that will shape the future of a field. Yet this process is under unprecedented strain, particularly in rapidly expanding areas such as artificial intelligence (AI), which face an explosion in paper submissions (Lipton and Steinhardt, 2019; Sculley et al., 2019; PaperCopilot, 2025). For the two largest AI conferences, submissions to the International Conference on Machine Learning (ICML) rose from $1676$ in 2017 to $12107$ in 2025, while those to the Conference on Neural Information Processing Systems (NeurIPS) increased from $3240$ in 2017 to $21575$ in 2025. This exponential growth has placed a heavy burden on peer review at AI conferences, leading to a reliance on graduate and undergraduate students as reviewers (Stelmakh et al., 2021), many without prior publications at these venues, and increasingly, the use of language models for review (Liang et al., 2024). This raises serious concerns about review quality. A striking example is the NeurIPS 2021 experiment, which revealed that about half of the accepted papers would have been rejected under a second independent review (Cortes and Lawrence, 2021; Beygelzimer et al., 2023). Such inconsistency underscores the difficulty of identifying research with lasting impact within the overwhelming volume of incremental work. A consequence of the decline of peer review that we cannot preclude is that the advancement of AI may follow a less-than-optimal path.
@@
Given the exhausted supply of qualified reviewers, we explore a vast, yet largely untapped, source of expert assessment: the authors themselves. Authors possess an unparalleled understanding of their own work, including its theoretical underpinnings, limitations, and the subtle details that are fundamental to its potential for long-term scientific impact (Su et al., 2025). This deep insight stands in contrast to the focus of many overburdened reviewers who, under tight deadlines, may prioritize quantifiable gains over a submission’s broader scientific implications. This is particularly true in AI research, where many reviewers are junior researchers who often focus on improvements in predictive accuracy (such as test accuracy on a benchmark dataset increasing from 91% to 92%) and may lack the experience to evaluate a paper’s long-term significance (Shah et al., 2018).
@@
To empirically test the hypothesis that author insight predicts scientific impact, a primary obstacle is soliciting candid self-assessments, as authors may be inclined to inflate the quality of their work. To overcome this, we designed and implemented a large-scale experiment at ICML in 2023, with the approval of the conference organizers. We asked authors who had submitted multiple papers to rank them according to their perceived quality and significance. This design elicits comparative rather than absolute judgments, meaning authors cannot simply claim all their submissions are of the highest quality. Under certain conditions, this comparative design incentivizes authors to truthfully report their true rankings (Su, 2025). The experiment yielded a rich dataset of self-rankings from 1,342 AI researchers, covering 2,592 submissions to ICML 2023, alongside their official review scores and acceptance decisions.
@@
Our analysis revealed that authors’ self-rankings are a powerful predictor of a paper’s future impact, measured by citations accumulated over 16 months. Citation count is a widely used metric for the scientific impact of research work (Aksnes et al., 2019). Submissions ranked highest by their authors received, on average, twice as many citations as those they ranked lowest, a trend that held for both accepted and rejected submissions. The predictive power was particularly striking for identifying high-impact work: of the 22 papers in our dataset that garnered over 150 citations, 18 were ranked highest by their authors. For comparison, we demonstrated that these self-rankings are a more accurate predictor of future citations than the review scores. These findings are highly statistically significant and remained robust after controlling for potential confounding factors such as early dissemination on preprint servers and authors’ self-citations. Our results suggest that authors’ self-rankings are a valuable, cost-effective signal that could be integrated into future review systems to help identify and elevate research of the highest impact.
1.1 Related work
A growing body of work explores mechanisms to improve peer review from the authors’ side (Aziz et al., 2019; Mattei et al., 2020; Srinivasan and Morgenstern, 2021). For example, Gardner et al. (2012) asked authors to rate their submissions to the Australasian Association for Engineering Education Annual Conference, and Rastogi et al. (2024) surveyed NeurIPS 2021 authors to estimate acceptance probabilities and provide pairwise comparisons of their submissions. Recently, a new approach to improving peer review for AI conferences starts by eliciting a ranking of submissions by the same author and uses the ranking to calibrate the (average) review scores (Su, 2021, 2025), [^1] A paper submitted to an AI conference is typically reviewed by multiple reviewers, and each reviewer assigns a score to the paper to evaluate its quality. with several extensions (Yan et al., 2025; Wu et al., 2023). This approach targets AI conferences where an author often submits multiple submissions. An empirical study showed that the calibrated scores were more accurate than review scores, but did not examine how these new scores related to the scientific impact of submissions (Su et al., 2025).
Many metrics are commonly used to capture paper quality or impact. Citation count, in particular, plays a central role in research evaluation, hiring, and funding decisions (Cole and Cole, 1971; Wang et al., 2013; Caon et al., 2020; Langfeldt et al., 2021). Prior work has also shown that citation patterns vary across paper types and topics (Miranda and Garcia-Carpintero, 2018; Small, 2018; Mendoza, 2021). In particular, Thelwall et al. (2023) found a particularly strong positive correlation between citation counts and paper quality in mathematical science and computer science papers compared to other fields. In Aksnes et al. (2019), the authors argued that citations can serve as indicators of scientific impact and relevance. Moreover, the relationship between final decisions, review scores, and citations were examined (Xie et al., 2024; Weitzner et al., 2024), and the associations between GitHub stars and citation counts were also analyzed (Zhou et al., 2024; Alrashedy and Binjahlan, 2024).
@@ 1.1 Related work
Figure 1: (a) The interface for the Phase One survey experiment, with summary statistics (see Section 2 for details and Figure S.1 for zoom-in view). (b) Comparison of citation counts between high- and low-ranked submissions, revealing a statistically significant difference ($P=1.05\times 10^{-6}$; see Section 3). (c) Proportions of high- and low-ranked submissions, categorized by how well they are cited (i.e., well-cited vs. less-cited) (see Section 4). (d) Mean citation counts for submissions partitioned into high- and low-value groups according to three metrics: self-rankings, pre-rebuttal scores, and post-rebuttal scores. A larger difference in means indicates greater predictive power for the metric (see Section 5).
@@ 2 Data collection
Our experiment was conducted in two phases. In Phase One, we carried out a survey-based experiment on OpenReview, which hosted the ICML 2023 peer review process for $6538$ submissions from $18535$ authors. The experiment itself was implemented with OpenRank.cc, a platform we developed for this purpose. On January 26, 2023, immediately after the submission deadline, an official email was distributed via OpenReview to all ICML authors requesting information about their submissions. Authors with multiple submissions were asked to rank their papers based on perceived quality, along with answering several questionnaire items. It is common for AI researchers to submit multiple papers to the same conference. Indeed, at ICML 2023, $4505$ of the $18535$ authors had more than one submission, and $5035$ of the total $6538$ submissions had at least one author having more than one submission. After the peer review process, we obtained review scores and final decisions for all submissions from OpenReview. In total, $5634$ authors completed the survey (response rate: $30.4\%$), of which $1342$ authors with multiple submissions provided rankings. Altogether, $2592$ submissions were ranked by at least one author, accounting for $39.6\%$ of all submissions.
@@
In Phase Two of the experiment, we used Semantic Scholar to collect citation data for each ranked submission. [^2] Semantic Scholar provides citation records for a paper within a specified time period. See Section B.3 for results based on Google Scholar. For each submission, we retrieved all papers that cited the submission between July 23, 2023 and November 22, 2024. For each citing paper, we extracted the title, author list, and publication date. A key challenge was accounting for multiple submission versions, which often involved revised titles or author lists before or after ICML 2023. Additionally, some titles or author names include special characters, such as “Schrödinger,” which introduced minor mismatches during automated title matching. To ensure accuracy, we retained only submissions with an exact match in both title and author list on Semantic Scholar (see Appendix A for details). After filtering, we obtained a final dataset consisting of 797 authors with multiple ranked submissions, each with valid citation data, involving a total of 1,527 unique submissions. All subsequent citation analyses are based on this filtered set. Across these submissions, the average number of citations per paper during the study window was $14.73$. Half of these submissions received at least $4$ citations, and one-quarter received at least $11$ citations. The most-cited submission received $979$ citations. [^3] These statistics are based on the $1527$ submissions with valid citation counts from Semantic Scholar.
@@ 4 Self-rankings identify well-cited papers on the edge
While we have shown that higher-ranked papers received substantially more citations on average, it is also important to assess predictivity in the edge case of highly cited papers. Building on the earlier observation, we next examined how well-cited submissions are distributed across authors’ self-rankings. As shown in Figure 1(c), among all submissions that received more than $150$ citations, $77.27\%$ were ranked first by at least one of their authors. Specifically, of the 22 submissions exceeding this threshold, 18 were ranked first, 3 were ranked second, and 4 were ranked last by at least one author. [^5] Only one submission was ranked both first and last by different authors, and one submission was neither ranked first nor last. Moreover, most well-cited papers belong to the high-ranked category.Focusing on high-ranked papers alone suffices to identify most well-cited papers.
@@
Figure 3: Most of well-cited papers fall into the high-ranked category. Empirical complementary cumulative distribution of citation counts, showing the proportion of papers with more than $C$ citations. The left panel presents accepted submissions; the right panel shows rejected ones. At every citation threshold $C$, a higher proportion of high-ranked papers exceed $C$ than low-ranked papers. Average citation counts for high- and low-ranked submissions are highlighted using vertical dashed lines.
@@ 5 Self-rankings outperform review scores in predicting citations
[…]
@@
Mean High 27.36 25.58 23.62 11.94 9.81 9.80
Low 13.82 17.64 18.18 6.32 8.82 8.38
$p$-value $9.34\times 10^{-3}$ $0.13$ $0.30$ $1.3\times 10^{-3}$ $0.59$ $0.43$
@@
50th percentile High 8.0 8.0 7.0 3.0 2.0 2.0
Low 6.0 6.0 6.5 2.0 3.0 3.0
25th percentile High 3.0 3.0 3.0 1.0 1.0 1.0
Low 3.0 3.0 3.0 1.0 1.0 1.0
Table 1: Authors’ self-rankings better identify impactful papers. Average number of citations (from July 23, 2023 to November 22, 2024) for (1) high-ranked versus low-ranked papers, (2) high-scored versus low-scored paper in post-rebuttal phase, and (3) high-scored versus low-scored paper in pre-rebuttal phase, grouped by their final decisions in ICML 2023. At a significance level of $P=0.05$, only the high-ranked papers show a significantly higher number of citations compared to the low-ranked papers, based on a paired $t$-test. In contrast, the differences in citation counts between high and low groups based on pre-rebuttal and post-rebuttal scores are not statistically significant.
@@
Figure 4 further validated the findings by stratifying submission pairs into four groups with similar pre- or post-rebuttal scores. For example, if an author ranked submission $A$ as better than $B$, this pair was only included if their review scores were close, for example, both between 5 and 6. Figure S.3 in the Supplementary Material shows the number of authors with multiple ranked submissions within each review score group. Even within groups of similar review scores, high-ranked papers by authors tended to accumulate more citations than their low-ranked counterparts. Notably, submissions with pre-rebuttal scores in the 5 to 6 range received the highest citation counts on average, driven by the two most-cited papers in our dataset, both ranked first by their authors. Moreover, the citation gap between high- and low-ranked papers widened as review scores increased. This suggested that while review scores correlate with paper impact, they may not be able to identify the most impactful work. In particular, in high-score regimes, authors’ self-rankings provided more complementary information that helped capture high-impact papers.
@@
Table 2: Distribution of 22 submissions with more than 150 citations. The largest proportion falls into the high-ranked group, while the smallest proportion falls into the low-ranked group, suggesting that authors’ self-rankings are better at identifying well-cited papers than review scores or final decisions. Note that, using the pre-rebuttal score column as an example, if a submission received the highest pre-rebuttal score among all of an author’s submissions, it is categorized into the high-scored group. The same applies for the low-scored group. Some papers with more than 150 citations may fall into both groups, or into neither.
As in Section 3, in addition, we placed the highest-scored submissions in one group and the lowest-scored submissions in the other. We first observed that authors’ self-rankings were more effective than either review scores or final decisions in identifying well-cited papers. In Table 2, among the 22 submissions with citation counts exceeding 150, 14 appeared in the high pre-rebuttal score group, compared to 18 in the high-ranked group. In contrast, 6 of these well-cited submissions fell into the low pre-rebuttal score group, compared to only 4 in the low-ranked group. Turning to final decisions, 15 of these submissions were accepted by ICML 2023, while 7 were rejected. Notably, 4 of the rejected submissions were nonetheless ranked as the top paper by at least one of their authors.
@@ 6 Control for the confounding factors
Another factor is authors’ preferences. First, authors may self-cite high-ranked papers more frequently. To test this, we calculate citation counts excluding self-citations, which are defined as citations from papers with no overlapping authors. Among 649 authors with multiple ranked submissions and valid non-self-citation counts, the middle and lower panels of Figure S.4 present cumulative averages from July 23, 2023 to November 22, 2024, conditional on final decisions and review scores. The same conclusion holds: high-ranked papers continue to receive more citations even when self-citations are excluded.

### v2

# How to Find Fantastic AI Papers: Self-Rankings as a Powerful Predictor of Scientific Impact Beyond Peer Review

@@ 1 Introduction
A central objective of peer review is to identify and promote research of high impact that will shape the future of a field. Yet, this process is under unprecedented strain, particularly in rapidly expanding areas such as artificial intelligence (AI), which face an explosion in paper submissions (lipton2019troubling; sculley2018avoiding; papercopilot2025statsicml). For the two largest AI conferences, submissions to the International Conference on Machine Learning (ICML) rose from $1676$ in 2017 to $12107$ in 2025, while those to the Conference on Neural Information Processing Systems (NeurIPS) increased from $3240$ in 2017 to $21575$ in 2025. This exponential growth has placed a heavy burden on peer review at AI conferences, leading to reliance on graduate and undergraduate students as reviewers (stelmakh2021novice), many without prior publications at these venues, and increasingly, the use of language models for review (liang2024can). This raises serious concerns about the quality of the review. A striking example is the NeurIPS 2021 experiment, which revealed that approximately half of the accepted papers would have been rejected under a second independent review (cortes2021inconsistency; beygelzimer2023has). This inconsistency highlights the challenge of identifying research with lasting impact amidst the overwhelming volume of incremental work. A consequence of the decline of peer review that we cannot preclude is that the advancement of AI may follow a less-than-optimal path.
@@
While the supply of qualified reviewers is strained, far less research has focused on eliciting expert assessment from the authors themselves (aziz2019strategyproof; mattei2020peernomination; srinivasan2021auctions; su2021you; rastogi2022authors). Authors possess an unparalleled understanding of their own work, encompassing its theoretical foundations, limitations, and the subtle details fundamental to its long-term scientific impact. This deep insight stands in contrast to the focus of many overburdened reviewers who, under tight deadlines, may prioritize quantifiable gains over a submission’s broader scientific implications. This is particularly true in AI research, where many reviewers are junior researchers who often focus on marginal improvements in predictive accuracy (such as a benchmark test accuracy increasing from 91% to 92%) and may lack the experience to evaluate a paper’s long-term significance (shah2018design).
@@
To empirically test the hypothesis that author insight predicts scientific impact, a primary obstacle is soliciting candid self-assessments, as authors may be inclined to inflate the quality of their work. One approach is to elicit a self-ranking of submissions from the same author, which is practical as it is common for the same author to submit multiple papers to the same conference in AI research. This design elicits comparative rather than absolute judgments, so authors cannot simply claim all their submissions are of the highest quality. Under certain conditions, this comparative design incentivizes authors to truthfully report their true rankings (su2021you; su2025you; yan2023isotonic). To empirically quantify the predictive power of authors’ self-rankings, we designed and implemented a large-scale experiment at ICML in 2023, with the approval of the conference organizers. We asked authors who had submitted multiple papers to rank them according to their perceived quality and significance. The experiment yielded a rich dataset of self-rankings from 1,342 AI researchers, encompassing 2,592 submissions to ICML 2023, along with their official review scores and acceptance decisions.
@@
Our analysis of the ICML 2023 data reveals that authors’ self-rankings are a powerful predictor of a paper’s future impact, measured by citations accumulated over 16 months. Citation count is a widely used, albeit imperfect, metric for the scientific impact of research (aksnes2019citations; huang2022evaluating). Submissions ranked highest by their authors received, on average, twice as many citations as those they ranked lowest, a trend that held for both accepted and rejected submissions, suggesting that self-rankings can not only denoise review scores but also capture a distinct dimension of impact (su2025icml). The predictive power was particularly striking for identifying high-impact work: of the 22 papers in our dataset that garnered over 150 citations, 17 were ranked highest by their authors. For comparison, we demonstrate that these self-rankings are a more accurate predictor of future citations than the review scores. These findings are highly statistically significant and remained robust after controlling for potential confounding factors such as early dissemination on preprint servers and authors’ self-citations. Our results suggest that authors’ self-rankings are a valuable, cost-effective signal that could be integrated into future review systems to help identify and elevate research of the highest impact.
@@
Figure 1: (a) The illustration for the Phase One survey experiment, with summary statistics (see Section 2 for details and Figure S.1 for the real interface). (b) Comparison of citation counts between high- and low-ranked submissions, revealing a statistically significant difference ($P=1.05\times 10^{-6}$; see Section 3). (c) Proportions of high- and low-ranked submissions, categorized by how well they are cited (i.e., well-cited vs. less-cited) (see Section 4). (d) Mean citation counts for submissions partitioned into high- and low-value groups according to three metrics: self-rankings, pre-rebuttal scores, and post-rebuttal scores. A larger difference in means indicates greater predictive power for the metric (see Section 5).
@@ 2 Data collection
Our experiment was conducted in two phases. In Phase One, we carried out a survey-based experiment on OpenReview, which hosted the ICML 2023 peer review process for $6538$ submissions from $18535$ authors. The experiment itself was implemented with OpenRank.cc, a platform we developed for this purpose. On January 26, 2023, immediately after the submission deadline, an official email was distributed via OpenReview to all ICML authors requesting information about their submissions. Authors with multiple submissions were asked to rank their papers based on perceived quality, along with answering several questionnaire items. In total, $5634$ authors completed the survey (response rate: $30.4\%$), of which $1342$ authors with multiple submissions provided rankings. Altogether, $2592$ submissions were ranked by at least one author, accounting for $39.6\%$ of all submissions. It is common for AI researchers to submit multiple papers to the same conference. Indeed, at ICML 2023, $4505$ of the $18535$ authors had more than one submission, and $5035$ of the total $6538$ submissions had at least one author having more than one submission. A paper submitted to an AI conference is evaluated by multiple reviewers. Before the rebuttal phase, every reviewer assigns a pre-rebuttal review score—at ICML 2023, on a 1–10 scale. During the rebuttal period, authors may respond to reviewers’ comments and clarify points without making substantial changes to the original submission. After considering the rebuttal, reviewers may revise their scores; these revised scores are referred to as post-rebuttal scores. We obtained the pre- and post-rebuttal scores, as well as the final decisions for all submissions, from OpenReview.
@@
In Phase Two of the experiment, we used Semantic Scholar to collect citation data for each ranked submission. [^1] Semantic Scholar provides citation records for a paper within a specified time period. See Section B.3 for results based on Google Scholar. For each submission, we retrieved all papers that cited the submission between July 23, 2023 and November 22, 2024. For each citing paper, we extracted the title, author list, and publication date. A key challenge was accounting for multiple submission versions, which often involved revised titles or author lists before or after ICML 2023. Additionally, some titles or author names include special characters, such as “Schrödinger,” which introduced minor mismatches during automated title matching. To ensure accuracy, we retained only submissions with an exact match in both title and author list on Semantic Scholar (see Section A for details). After filtering, we obtained a final dataset consisting of 797 authors with multiple ranked submissions, each with valid citation data, involving a total of 1,527 unique submissions. All subsequent citation analyses are based on this filtered set. Across these submissions, the average number of citations per paper during the study window was $14.73$. Half of these submissions received at least $4$ citations, and one-quarter received at least $11$ citations. The most-cited submission received $979$ citations. [^2] These statistics are based on the $1527$ submissions with valid citation counts from Semantic Scholar.
@@ 4 Self-rankings identify well-cited papers on the edge
While we have shown that higher-ranked papers received substantially more citations on average, it is also important to assess predictivity in the edge case of highly cited papers. Building on the earlier observation, we next examined how well-cited submissions are distributed across authors’ self-rankings. As shown in Figure 1(c), among all submissions that received more than $150$ citations, $77.27\%$ were ranked first by at least one of their authors. Specifically, of the 22 submissions exceeding this threshold, 17 were ranked first, 3 were ranked second, and 5 were ranked last by at least one author. Only one submission was ranked both first and last by different authors, while another was neither ranked first nor last. We chose the top 22 submissions because $22/1527\approx 1.44\%$, which roughly corresponds to the proportion of “Accepted as Oral” at major AI conference. Moreover, most well-cited papers belong to the high-ranked category. Focusing on high-ranked papers alone suffices to identify most well-cited papers.
@@
Figure 3: Most of the well-cited papers fall into the high-ranked category. Empirical complementary cumulative distribution of citation counts, showing the proportion of papers with more than $C$ citations. The left panel presents accepted submissions; the right panel shows rejected ones. At every citation threshold $C$, a higher proportion of high-ranked papers exceed $C$ than low-ranked papers. Average citation counts for high- and low-ranked submissions are highlighted using vertical dashed lines.
@@ 5 Self-rankings outperform review scores in predicting citations
We further compare the predictive power of authors’ self-rankings versus reviewer scores for submissions’ future citation counts. Conditioning on the final decision, we analyze 280 authors with multiple accepted and ranked submissions and 343 authors with multiple rejected and ranked submissions. As shown in Table 1 and Figure 1(d), the citation-count gap between high and low category is largest when submissions are grouped by self-rankings, compared with reviewer scores, both pre- and post-rebuttal. In particular, only authors’ self-rankings yield statistically significant differences in citation counts between high- and low-ranked papers (paired $t$-test, $P=9.34\times 10^{-3}$). This indicates that self-rankings are the most predictive of submissions’ citation counts. These patterns are even more particularly pronounced among rejected submissions, possibly because reviewers may not differentiate scores as carefully once a paper falls below the acceptance threshold. Notably, the most-cited paper in our dataset, which accumulated 979 citations between July 23, 2023 and November 22, 2024, was ranked first by its authors but received a post-rebuttal score of 5, lower than that of their other submissions, and was therefore accepted only as a poster.
@@
Mean High 27.36 25.58 23.62 12.41 9.99 9.98
Low 13.82 17.64 18.18 6.36 9.15 8.70
$p$-value $9.34\times 10^{-3}$ $0.13$ $0.30$ $7.7\times 10^{-4}$ $0.65$ $0.49$
@@
Table 1: Authors’ self-rankings better identify impactful papers. Average number of citations (from July 23, 2023 to November 22, 2024) for (1) high-ranked versus low-ranked papers, (2) high-scored versus low-scored papers in the post-rebuttal phase, and (3) high-scored versus low-scored papers in the pre-rebuttal phase, grouped by their final decisions in ICML 2023. At a significance level of $P=0.05$, only the high-ranked papers show a significantly higher number of citations compared to the low-ranked papers, based on a paired $t$-test. In contrast, the differences in citation counts between high and low groups based on pre-rebuttal and post-rebuttal scores are not statistically significant.
@@
Figure 4 further validates the findings by stratifying submission pairs into four groups with similar pre- or post-rebuttal scores. For example, an author ranks a submission $A$ as better than $B$, and this pair will only be included if their review scores are close, for example, both between 5 and 6. Figure S.3 in the Supplementary Material shows the number of authors with multiple ranked submissions within each review score group. Even within groups of similar review scores, high-ranked papers by authors tended to accumulate more citations than their low-ranked counterparts. Notably, submissions with pre-rebuttal scores in the 5 to 6 range received the highest citation counts on average, driven by the two most-cited papers in our dataset, both ranked first by their authors. Moreover, the citation gap between high- and low-ranked papers widens as review scores increase. This suggests that while review scores correlates with paper impact, they may not be able to identify the most impactful work. In particular, in high-score regimes, authors’ self-rankings provide more complementary information that helps capture high-impact papers.
@@
Table 2: Distribution of 22 submissions with more than 150 citations. The largest proportion falls into the high-ranked group, while the smallest proportion falls into the low-ranked group, suggesting that authors’ self-rankings are better at identifying well-cited papers than review scores or final decisions. Note that among 22 submissions that received more than 150 citations, 14 having highest pre-rabuttal scores among all of an author’s submissions, 6 having lowest pre-rebuttal scores. One submission having both highest and lowest pre-rebuttal scores by different authors and 3 submissions having neither highest or lowest pre-rebuttal scores among all of their authors.
As in Section 3, in addition, we place the highest-scored submissions in one group and the lowest-scored submissions in the other. We first observe that authors’ self-rankings are more effective than either review scores or final decisions in identifying well-cited papers. In Table 2, among the 22 submissions with citation counts exceeding 150, 14 appeared in the high pre-rebuttal score group, compared to 17 in the high-ranked group. In contrast, 6 of these well-cited submissions fell into the low pre-rebuttal score group, compared to 5 in the low-ranked group. Turning to final decisions, 15 of these submissions were accepted by ICML 2023, while 7 were rejected. Notably, 4 of the rejected submissions were nonetheless ranked as the top paper by at least one of their authors.
@@ 6 Control for the confounding factors
Another factor is the authors’ preferences. First, authors may self-cite high-ranked papers more frequently. To test this, we calculate citation counts excluding self-citations, which are defined as citations from papers with no overlapping authors. Among 649 authors with multiple ranked submissions and valid non-self-citation counts, the middle and lower panels of Figure S.4 present cumulative averages from July 23, 2023 to November 22, 2024, conditional on final decisions and review scores. The same conclusion holds: high-ranked papers continue to receive more citations even when self-citations are excluded.